SPECTROSCOPIC DETECTION OF THE HYDROTRIOXYL RADICAL (HOOO) IN THE GAS PHASE

\underline{KOHSUKE SUMA}, YOSHIHIRO SUMIYOSHI, AND YASUKI ENDO, Department of Basic Sciences, Graduate School of Arts and Sciences, The University of Tokyo, Komaba, Meguro-ku, Tokyo 153-8902, Japan.

Pure rotational spectra of the HOOO and DOOO radicals have been observed in a supersonic jet by using a Fourier transform microwave (FTMW) and FTMW-millimeter wave double resonance spectroscopy with a pulsed discharge nozzle. From a total of 12 observed rotational transitions with fine and hyperfine splittings, precise molecular constants have been determined. The experimentally determined molecular structure is trans planar, which was not expected before the present experimental detection, since the large number of ab initio calculations so far reported have predicted its structure to be cis planar. The determined bond length between the HO and O_{2} moieties, HO-OO, is fairly long, 1.688 Å, and the molecular structure of HOOO is similar to that of an isoelectronic radical, FOO, which also has a fairly long FO bond.